A 1.50 V battery supplies 0.200 W of power to a small flashlight for 25.0 min. (a) How much charge (in C) does it move? (b) How many electrons must move to carry this charge? Additional Materials G Reading 2. -11 points 0SColPhu2016 19.1.XP 001. 0/20 Submissions Used when removing a wool sweater, a static discharge of 1.00 pc dissipates 4.00 x 10-3 J of energy. What voltage (in V) was involved? Additional Materials Reading

Please complete both thank you
Answer #1

1.

W = V x I
I = W/V
I = 0.200/1.50
I, the current = 0.13333A

(a) Q = I x t
Q = 0.13333 x 25 x 60
Q = 199.9C or Q=200C

(b) Electron has charge, e= -1.6 x 10-19 C
No of electrons = Q/e = 200/1.6 x 10-19 = 1.25 x 1021

Post another query in separate thread